### v1

# On concentric fractal spheres and spiral shells

Abstract.
We investigate dimension theoretic properties of concentric topological spheres, which are fractal sets emerging both in pure and applied mathematics. We calculate the box dimension and Assouad spectrum of such collections, and use them to prove that fractal spheres cannot be shrunk into a point at a polynomial rate. We also apply these dimension estimates to quasiconformally classify certain spiral shells, a generalization of planar spirals in higher dimensions. This classification also provides a bi-Hölder map between shells and constitutes an addition to a general programme of research proposed by J. Fraser in [23].
Keywords. Fractal, dynamical systems, spiral, Assouad spectrum, quasiconformal mappings.
2020 Mathematics Subject Classification:
Primary 28A80; Secondary 37C45, 54A20, 30C65
1. Introduction
Concentric objects occur in various areas of pure and applied mathematics. In complex dynamics, the Fatou components of entire transcendental maps are typically of fractal boundary and converge to infinity (see for instance [3, 8]). Therefore, under a conformal inversion that keeps most geometric properties intact, the components of the map can be seen as fractals “centered” at 00. In the context of harmonic analysis, concentric Euclidean spheres are closely associated to weights of measures that lie in the important Muckenhoupt class [2]. In fact, polynomially-concentric sphere collections, i.e., unions of spheres $S_{n}$ centered at 00 with radii equal to $n^{-p}$, for some $p>0$, provide one of the very few examples of bounded sets whose distance function lies in the Muckenhoupt class, for certain exponents.
A set $S\subset\mathbb{R}^{d}$ is called a topological sphere, if there is a homeomorphism $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ which maps the unit Euclidean sphere centered at 00, denoted by $S(0,1)\subset\mathbb{R}^{d}$, onto $S$. A topological sphere can potentially be smooth, or extremely “rough”. A quantitative classification between the two cases is a difficult problem of many interpretations. As a result, it has received attention within numerous fields, such as topology, analysis on metric spaces and geometric group theory (see [9] for a detailed survey on the topic). Despite such difficulties on the theoretical front, concentric topological spheres provide examples of concentric fractals that emerge in various applied areas, such as mathematical physics [13], [32], machine learning [21], and dynamic scaling of mathematical models [11].
It might seem natural that very rough fractals, which are disjoint and centered at a fixed point, would have to decrease a lot in size while they approach the center. More specifically, in concentric topological sphere collections, where the spheres are allowed to be fractals of high fractal dimension, they need to converge to their center significantly fast in order to avoid overlaps. Hence, for any concentric collection of topological spheres, there are certain restrictions on either their dimension, or on how fast they accumulate around their center. One of our main results is addressing this phenomenon for fractal spheres, by studying certain dimension notions for the whole collection. For what follows, we fix an integer $d\geq 2$, and we denote by $B(0,1)\subset\mathbb{R}^{d}$ the open unit ball in $\mathbb{R}^{d}$ centered at 00. We formalize the notion of concentric sphere collections in the following definition.
Figure 1. An example of a concentric collection of snowflakes.
Definition 1.1.
Let $S_{0}\subset\mathbb{R}^{d}$ be a topological sphere with $S_{0}=f(S(0,1))$, for some homeomorphism $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. For $x_{0}\in f(B(0,1))$ and a sequence of similarities $\{g_{n}\}_{n\in\mathbb{N}}$, set $S_{n}:=g_{n}(S_{0})$. We say that the union $S:=\bigcup_{n\in\mathbb{N}_{0}}S_{n}$ is a concentric sphere collection centered at $x_{0}$ and generated by $S_{0}$, if there are $c_{1},c_{2}\in(0,1]$, $n_{0}\in\mathbb{N}$, and a strictly decreasing sequence $\{a_{n}\}_{n\in\mathbb{N}}$ converging to 00 such that:
  1. (i)
    $d_{H}(S_{n},\{x_{0}\})=c_{1}a_{n}$, and
  2. (ii)
    $\operatorname{dist}(S_{n},S_{n+1})=c_{2}(a_{n}-a_{n+1}),$
for all $n\geq n_{0}$. Moreover, if there is $p>0$ so that $a_{n}=n^{-p}$, for all $n\geq n_{0}$, then we say that $S$ is a $p$-(polynomially) concentric sphere collection centered at $x_{0}$ and generated by $S_{0}$, and we denote it by $S_{p}=S_{p}(x_{0},S_{0})$.
An example of a planar concentric sphere collection is depicted in Figure 1, where $S_{0}$ is the Koch snowflake (see [18]). Note that in Definition 1.1, $\operatorname{dist}(S_{n},S_{n+1})$ denotes the usual Euclidean distance between the sets $S_{n},S_{n+1}$, and $d_{H}$ denotes the Hausdorff distance (see Section 2 for the definition). The latter is a natural way to talk about convergence of a collection of sets into a point (see for instance [34, Section 4.13], [22, Chapter 5]). Therefore, one could perceive Definition 1.1 as the process of “shrinking” the topological sphere $S_{0}$ into a point $x_{0}$, using contracting similarities $g_{n}$ of ratio approximately $a_{n}$. Such problems have attracted a lot of interest from various areas of mathematics, and they have been extensively studied both from a deterministic [6, 7] and a probabilistic [40] point of view. One of the main contributions of this paper is to emphasize that dimension-theoretic techniques can be crucial to the understanding of this phenomenon. Namely, we adopt such techniques in order to prove that a topological sphere $S_{0}$ of non-trivial box dimension $\dim_{B}S_{0}$ (see Section 2 for the definition) cannot give rise to polynomially concentric sphere collections of any degree.
Theorem 1.2.
Let $S_{0}\subset\mathbb{R}^{d}$ be a topological sphere with $\dim_{B}S_{0}\in(d-1,d]$. For any $p>0$, there is no $p$-concentric sphere collection generated by $S_{0}$.
It should be noted that it is always possible to shrink a topological sphere $S_{0}$ into a point through a sequence of similarities as described in Definition 1.1, as long as the sequence $a_{n}$ converges fast enough to 00. See Section 5 for a detailed discussion. Theorem 1.2 establishes that sequences of the form $n^{-p}$ (or of slower rate of convergence to 00) are simply not fast enough to shrink a topological sphere of non-trivial box dimension. Hence, all concentric fractal spheres encountered in the aforementioned applications [13, 32, 21, 11] either “smoothen” out as they approach the common center, or they shrink at an exponential rate. On the other hand, there are still various examples of topological spheres in $\mathbb{R}^{d}$ of box dimension equal to $d-1$ that generate polynomially concentric collections. For instance, the boundary of any convex set homeomorphic to $B(0,1)\subset\mathbb{R}^{d}$ can be polynomially shrunk into its center of mass.
The proof of Theorem 1.2 is achieved by assuming towards contradiction that for $d_{0}>d-1$ and $p>0$, a $p$-concentric sphere collection $S_{p}(x_{0},S_{0})\subset\mathbb{R}^{d}$ exists. We then use dimension theoretic tools, and specifically the Assouad spectrum notion $\{\dim_{A}^{\theta}S_{p}:\theta\in(0,1)\}$ introduced by Fraser-Yu [24] (see Section 2 for the definition), in order to show that such a collection would be of dimension larger than $d$, surpassing the dimension of the ambient space $\mathbb{R}^{d}$ and leading to a contradiction. A partial version of Theorem 1.2 can be proved by calculating the box dimension of such a collection, which results in a contradiction only for certain values of $p$. Hence, the Assouad spectrum is essential in achieving the result in full generality with this approach. Theorem 1.2 is the first application of the Assouad spectrum in this context of “rigidity” of fractal collections. Moreover, the property of the Assouad spectrum to trace “finer” geometric data is also crucial for “classification” problems under families of mappings. We establish Assouad spectrum estimates for concentric sphere collections (see Section 3), which we apply to classify certain objects, up to classes of homeomorphisms.
An instance of a planar set resembling a collection of concentric circles is the spiral of the form
$\mathcal{S}(\phi):=\{\phi(t)e^{it}:\,t>1\},$
where $\phi:[1,\infty)\rightarrow(0,\infty)$ is a continuous, strictly decreasing function, with $\lim_{t\rightarrow\infty}\phi(t)=0$. Indeed, such spirals share many geometric properties with concentric collections of circles of the form $\cup_{n\in\mathbb{N}}S(0,\phi(n))$, while they hold a prominent role in fluid turbulence [19, 36, 43, 44], dynamical systems [45, 27], and even certain types of models in mathematical biology [37]. Moreover, they provide examples of “non-intuitive” fractal behavior (see [17]), while they have also been extensively studied due to their unexpected analytic properties. For instance, Katznelson-Nag-Sullivan [29] demonstrated the dual nature of such spirals, lying in-between smoothness and “roughness”, as well as their connection to certain Riemann mapping questions. The relation between the rate of convergence of $\phi(t)$ for $t\rightarrow\infty$ and the existence of Lipschitz and Hölder parametrizations has also been extensively studied by the aforementioned authors in [29], by Fish-Paunescu in [20], and by Fraser in [23]. In particular, Fraser in the latter paper focuses on spirals that resemble polynomially concentric circles, i.e., spirals where $\phi(t)=t^{-p}$, for $p>0$, and emphasizes further the “fractal” nature of these objects, due to the peculiar behavior they exhibit in relation to certain dimension notions.
A natural higher dimensional generalization of spirals is the notion of spiral shells in $\mathbb{R}^{3}$. They exhibit similar behavior to spirals, which has attracted the interest of researchers in physics [38, 39] and mathematical biology [37]. For $p>0$, we consider polynomial (spiral) shells of the form
$\mathcal{S}_{p}:=\{(u^{-p}\cos u\sin v,u^{-p}\sin u\sin v,u^{-p}\cos v)\in\mathbb{R}^{3}:\,u,\,v\in[1,\infty)\}\subset\mathbb{R}^{3}.$
The resemblance to the polynomial spirals mentioned above is perhaps already evident from the definition. In addition, spiral shells have a very similar geometric structure to $p$-polynomially concentric spheres in $\mathbb{R}^{3}$. See Figure 2.
Figure 2. An example of a polynomial spiral shell for $p=1/2$.
One significant restriction in the analytic and geometric study of these objects is the lack of planar tools in $\mathbb{R}^{3}$. For instance, Liouville’s theorem states that all conformal maps are Möbius transformations in $\mathbb{R}^{3}$ (see for instance [41, p. 15, 43]). This was one of the motivations for the definition of quasiconformal maps in $\mathbb{R}^{d}$, $d\geq 2$ (see Section 4 for formal definition). More specifically, while conformal maps transform infinitesimal balls into infinitesimal balls, quasiconformal maps transform them into infinitesimal ellipsoids of globally bounded eccentricity. The bound on said eccentricity is called dilatation. If the dilatation of a quasiconformal map $f$ is at most some uniform constant $K\geq 1$, we say that $f$ is a $K$-quasiconformal map (see Section 4). The theory of quasiconformal maps has a rich history, which still continues to develop actively, especially due to its many connections and applications within various fields of mathematics. We refer to the books by Lehto and Virtanen [31], Väisälä [41], and Heinonen [26] for thorough expositions on these maps and their importance in $\mathbb{C}$, $\mathbb{R}^{d}$ and metric spaces, respectively.
Distinguishing sets up to quasiconformal maps, particularly at the level of the dilatation, is in general a hard problem. For instance, determining under what assumptions a topological sphere is a quasiconformal image of the usual unit sphere $S(0,1)$ is a long-standing open problem, with many implications in geometric group theory and topology (see [9] for a survey on this topic and connections to Cannon’s conjecture, see also [10]). For planar topological spheres, i.e., simple closed curves homeomorphic to the unit circle, this classification problem was solved by Ahlfors in [1], and their dimension-theoretic properties were fully determined by Smirnov in [42]. In higher dimensions only partial progress has been made; see [10, 35] and the references therein.
Given the ties of spiral shells to several applications, as well as their relation to concentric topological spheres, it is natural to try and classify them under quasiconformal maps. Indeed, this is the outcome of our second main result. By applying the Assouad spectrum estimates we establish for concentric spheres (see Theorem 3.3), along with the dimension distortion properties of quasiconformal maps, proved by Tyson and the author in [16], we achieve the desired classification of polynomial shells.
Theorem 1.3.
For $p\geq q>0$, let $\mathcal{S}_{p}$ and $\mathcal{S}_{q}$ be two polynomial spiral shells. There is a $K$-quasiconformal mapping $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with $F(S_{p})=S_{q}$ if, and only if, $K\geq p/q$.
Note that quasiconformal maps are also locally Hölder continuous, with locally Hölder continuous inverse (see Section 4 for definitions). Therefore, in the proof of Theorem 1.3, the quasiconformal map we construct, mapping $\mathcal{S}_{p}$ onto $\mathcal{S}_{q}$, is a bi-Hölder map between polynomial spiral shells, with the Hölder exponents quantitatively depending on the contracting ratios of the shells $p,q$. This contributes to a general direction proposed by J. Fraser in [23], on bi-Hölder classification of sets in $\mathbb{R}^{d}$. We discuss this connection further in Section 5.
This paper is organized as follows. Section 2 includes the notation we follow, and all the relevant dimension notions and their properties. In Section 3 we prove precise estimates for the box dimension and Assouad spectrum of $p$-concentric sphere collections, and apply similar dimension-theoretic arguments to prove Theorem 1.2. In Section 4 we recall the formal definition of quasiconformality, and state the relevant dimension distortion result we need for quasiconformal maps. Moreover, we demonstrate the shared dimension-theoretic properties between polynomially concentric spheres and spiral shells, which allows us to apply the quasiconformal dimension distortion result in order to classify polynomial shells as stated in Theorem 1.3. Section 5 contains further remarks and potential future directions that relate to this work.
Acknowledgements
The author wishes to thank Vasiliki Evdoridou, Manisha Garg and Carlos Mudarra for the interesting and motivating discussions on this work.
2. Background
We first establish the notation we follow. Given two positive functions $h_{1},h_{2}:(0,\infty)\rightarrow(0,\infty),$ we write $h_{1}(r)\lesssim h_{2}(r)$ if there is $C>0$, independent of $r$, such that $h_{1}(r)\leq Ch_{2}(r)$, for all $r>0$. We call $C>0$ the comparability constant of the relation $h_{1}(r)\lesssim h_{2}(r)$. We write $h_{1}(r)\gtrsim h_{2}(r)$ if $h_{2}(r)\lesssim h_{1}(r)$. Lastly, we write $h_{1}(r)\simeq h_{2}(r)$ if both $h_{1}(r)\lesssim h_{2}(r)$ and $h_{1}(r)\gtrsim h_{2}(r)$ hold. Note that $h_{1}(r)\simeq h_{2}(r)$ in general implies that there are $C_{1},C_{2}>0$, potentially distinct, such that $C_{1}h_{1}(r)\leq h_{2}(r)\leq C_{2}h_{1}(r)$ for all $r>0$. However, this implies that there is a single $C=\max\{C_{1}^{-1},C_{2}\}>0$ such that $C^{-1}h_{1}(r)\leq h_{2}(r)\leq Ch_{1}(r)$ for all $r>0$. In this case, we say that $C$ is the comparability constant of the relation $h_{1}(r)\simeq h_{2}(r)$. Given $x\in\mathbb{R}^{d}$, $r>0$, we denote by $B(x,r)$ the open ball centered at $x$ of radius $r$, and by $S(x,r)=\partial B(x,r)$ the corresponding sphere in $\mathbb{R}^{d}$. Recall that a set $S\subset\mathbb{R}^{d}$ is called a topological sphere, if there is a homeomorphism $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with $S=f(S(0,1))$. Given two compact sets $A,B\subset\mathbb{R}^{d}$, we denote their usual distance by $\operatorname{dist}(A,B)$, and their Hausdorff distance by
$d_{H}(A,B)=\max\Bigl{\{}\sup\{\operatorname{dist}(\{a\},B):\,a\in A\},\,\sup\{\operatorname{dist}(A,\{b\}):\,b\in B\}\Bigr{\}}.$
We next define the dimension notions we need, along with their properties. We refer to the books by Falconer [18] and Fraser [22] for thorough expositions on these topics. Let $E\subset\mathbb{R}^{d}$ be a non-empty bounded set. For $r>0$, we denote by $N_{r}(E)$ the smallest number of $r$-cubes needed to cover $E$, i.e., cubes that lie in the axes-oriented $r$-mesh of $\mathbb{R}^{d}$ and intersect $E$ (see [22, p. 8]). The lower and upper box dimensions of $E$ are defined by
$\underline{\dim_{B}}E=\liminf_{r\to 0}\,\frac{\log N_{r}(E)}{-\log r}\qquad\text{and}\qquad\overline{\dim_{B}}E=\limsup_{r\to 0}\,\frac{\log N_{r}(E)}{-\log r},$
respectively. If the lower and upper box dimensions are equal, we call the common value the box dimension of $E$ and denote it by $\dim_{B}E$.
Note that we have $\dim_{B}E=d_{0}$ if, and only if, for all $\epsilon>0$ there is $r_{0}=r_{0}(\epsilon)\in(0,1)$ such that
(2.1) $r^{-d_{0}+\epsilon}\leq N_{r}(E)\leq r^{-d_{0}-\epsilon},$
for all $r<r_{0}$.
Recall that a map $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a similarity if $g(x)=c_{g}\mathbf{A}x+t$ for all $x\in\mathbb{R}^{d}$, where $c_{g}>0$, $\mathbf{A}\in\mathcal{O}(\mathbb{R},d)$ is a real orthogonal matrix, and $t\in\mathbb{R}^{d}$. We call the scalar $c_{g}$ the similarity ratio of $g$. A standard property of the covering number $N_{r}(E)$ that we need is the way it is distorted by similarities. More specifically, if $g$ is a similarity map with similarity ratio $c$, then
$N_{r}(g(E))=N_{r/c}(E),$
for all $r\in(0,1)$ (see, for instance, [18] Proposition 2.5). Given that $\dim_{B}E=d_{0}$, using the above, along with (2.1) results in
(2.2) $\left(\frac{r}{c}\right)^{-d_{0}+\epsilon}\leq N_{r}(E)\leq\left(\frac{r}{c}\right)^{-d_{0}-\epsilon},$
for all $r<\min\{cr_{0},r_{0}\}$. This gives a relation between covering numbers of similar topological spheres that is crucial in our arguments in Section 3.
For an arbitrary (not necessarily bounded) set $E\subset\mathbb{R}^{d}$, the Assouad dimension of $E$ is
$\dim_{A}E=\inf\left\{\alpha>0\,:\,{\exists\,C>0\mbox{ s.t. }N_{r}(B(x,R)\cap E)\leq C(R/r)^{\alpha}\atop\mbox{ for all $0<r\leq R$ and all $x\in E$}}\right\}.$
While we do not use the Assouad dimension definition directly, we include it for the sake of completeness, and to relate it to the notion of the Assouad spectrum.
The Assouad spectrum, introduced by Fraser and Yu [24], is a one-parameter family of metrically defined dimensions which interpolates between the upper box dimension and the Assouad dimension. Specifically, the Assouad spectrum of a set $E\subset\mathbb{R}^{d}$ is a collection of values $\{\dim_{A}^{\theta}E:0<\theta<1\}$, where
$\dim_{A}^{\theta}E=\inf\left\{\alpha>0\,:\,{\exists\,C>0\mbox{ s.t. }N_{r}(B(x,r^{\theta})\cap E)\leq C(r^{\theta}/r)^{\alpha}\atop\mbox{ for all $0<r<1$ and all $x\in E$}}\right\},$
for all $\theta\in(0,1)$. We may also refer to the value $\dim_{A}^{\theta}E$ as the $(\theta)$-Assouad spectrum of E, for a specific $\theta\in(0,1)$.
The Assouad spectrum essentially captures the growth rate of the covering number $N_{r}(B(x,R)\cap E)$ for scales $10<r\leq R<$ 1 related by $R=r^{\theta}$. The map $\theta\mapsto\dim_{A}^{\theta}E$ is continuous at all $10<\theta<$ 1 [22, Section 3.3.3], and for a fixed bounded set $E\subset\mathbb{R}^{d}$ we have
$\dim_{A}^{\theta}E\to{\overline{\dim_{B}}}E\quad\mbox{as $\theta\to 0$},\qquad\dim_{A}^{\theta}E\to\dim_{qA}E\quad\mbox{as $\theta\to 1$},$
where $\dim_{qA}E$ denotes the quasi-Assouad dimension of $E$, a variant of Assouad dimension introduced by Lü and Xi [33]. We always have $\dim_{qA}E\leq\dim_{A}E$, and equality holds in many situations, including all sets of interest in this paper. We refer to [22, Chapter 3] for proofs of the above properties.
In Section 4 we use a slightly modified version of the Assouad spectrum, where the relationship $R=r^{\theta}$ between the two scales of the Assouad dimension definition is relaxed to an inequality $R\geq r^{\theta}$. Namely, given $E\subset\mathbb{R}^{d}$ and $\theta\in(0,1)$, we set
$\dim_{A,reg}^{\theta}E:=\inf\left\{\alpha>0\,:\,{\exists\,C>0\mbox{ s.t. }N_{r}(B(x,R)\cap E)\leq C(R/r)^{\alpha}\atop\mbox{ for all $0<r\leq R^{1/\theta}<R$ and all $x\in E$}}\right\},$
This modification leads to the notion of upper, or regularized Assouad spectrum, denoted by $\{{\dim_{A,reg}^{\theta}}E:0<\theta<1\}$. See [24] and [22, Section 3.3.2] for more information. Tyson and the author focused on this notion in [16], as it aligns more naturally with properties of the Assouad dimension that are fundamental in dimension distortion results (see also [14]). For a fixed $\theta\in(0,1)$ the key relationship between the respective two values in the two spectra (see [22, Theorem 3.3.6]) is that
(2.3) ${\dim_{A,reg}^{\theta}}E=\sup_{0<\theta^{\prime}<\theta}\dim_{A}^{\theta^{\prime}}E.$
It is immediate from (2.3) that if $\dim_{A}^{\theta}E$ is increasing in $\theta$, then $\dim_{A,reg}^{\theta}E=\dim_{A}^{\theta}E$. This is particularly useful, since $\dim_{A}^{\theta}E$ is often easier to calculate, while the dimension distortion results we use in Section 4 are proved for the regularized Assouad spectrum.
Moreover, the relation between the different dimensions for a fixed bounded set $E$ follows by (2.3) and the respective definitions. Namely,
(2.4) ${\overline{\dim_{B}}}E\leq\dim_{A}^{\theta}E\leq\dim_{A,reg}^{\theta}E\leq\dim_{\mathrm{qA}}E\leq\dim_{A}E,$
for all $\theta\in(0,1)$. We also state an upper bound on the Assouad spectrum involving the box dimension, which reduces certain calculations in arguments within Section 3 (see [22, Lemma 3.4.4]).
Lemma 2.1.
Given a bounded non-empty set $E\subset\mathbb{R}^{d}$ with $d_{0}:=\dim_{B}E$, we have
$\dim_{A}^{\theta}E\leq\min\left\{\frac{d_{0}}{1-\theta},\,d\right\},$
for all $\theta\in(0,1)$.
3. Dimensions and impossible concentric spheres
In this section, we first establish certain dimension-theoretic properties for polynomially concentric spheres. Namely, we estimate the box dimension and Assouad spectrum of such collections. We then modify the arguments we employed while proving the aforementioned estimates, in order to show the restrictions of polynomially concentric sphere collections with regard to the generating sphere, which yields Theorem 1.2.
Let $p>0$ and $S_{0}=f(S(0,1))\subset\mathbb{R}^{d}$ be a topological sphere, for some homeomorphism $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, with $d_{0}:=\dim_{B}S_{0}\leq d$. Note that the topological dimension of $S(0,1)$ is equal to $d-1$, and, since the topological dimension is preserved under homeomorphisms, so is the topological dimension of $S_{0}$. Thus, we immediately have $d_{0}\geq d-1$ (see, for instance, [26] p. 62). Suppose that $S_{p}=\cup_{n=0}^{\infty}S_{n}$ is a $p$-concentric sphere collection centered at some $x_{0}\in\mathbb{R}^{d}$ generated by $S_{0}$, as in Definition 1.1.
We can perform certain reductions on $S_{p}$. Without loss of generality, we can assume that $x_{0}=0$, since all the dimension notions we consider are translation-invariant. In addition, by finite stability of the box dimension and the Assouad spectrum (see [22], p. 18, 49), we can further assume that for the given $S_{p}$, we have $n_{0}=1$ in Definition 1.1. The above reductions imply that there is $c_{1}\in(0,1]$ such that
(3.1) $d_{H}(S_{n},\{0\})=c_{1}n^{-p},$
for all $n\in\mathbb{N}$. In addition, there is $c_{2}\in(0,1]$, independent of $n$, such that
(3.2) $\operatorname{dist}(S_{n},S_{n+1})=c_{2}(n^{-p}-(n+1)^{-p}),$
for all $n\in\mathbb{N}$. Moreover, for any $n\in\mathbb{N}$, we have that
(3.3) $\operatorname{dist}(S_{n},\{0\})\simeq n^{-p},$
where the comparability constant is independent of $n$. Indeed, the upper bound for (3.3) follows trivially from (3.1). In order to prove that $\operatorname{dist}(S_{n},\{0\})\gtrsim n^{-p}$, let $z_{n}\in S_{n}$ be a point such that $|z_{n}|=\operatorname{dist}(S_{n},\{0\})$. Since $S_{j}$ for $j\in\{n+1,n+2,\dots\}$ are all topological spheres with corresponding topological balls containing 00, the line segment with end-points at $z_{n}$ and 00 intersects each such $S_{j}$, say at $z_{j}\in S_{j}$. This implies that
$|z_{j}-z_{j+1}|\geq\operatorname{dist}(S_{j},S_{j+1})\simeq j^{-p}-(j+1)^{-p},$
for all $j\in\{n,n+1,\dots\}$. However, the sum of all of these distances $|z_{j}-z_{j+1}|$ has length less than the line segment where all these points lie on, namely
$\operatorname{dist}(S_{n},\{0\})\gtrsim\sum_{j=n}^{\infty}(j^{-p}-(j+1)^{-p})=\sum_{j=1}^{\infty}(j^{-p}-(j+1)^{-p})-\sum_{j=1}^{n-1}(j^{-p}-(j+1)^{-p}).$
Using the fact that for any $m\in\mathbb{N}$ we have $\sum_{j=1}^{m}(j^{-p}-(j+1)^{-p})=1-(1+m)^{-p}$, the proof of (3.3) is complete. Note that (3.3) also implies that the similarity ratio of $g_{n}$ is in fact comparable to $n^{-p}$, for all $n\in\mathbb{N}$. This allows us to use (2.2) in our arguments.
The above reductions are enough to start with the calculation of the box dimension of $S_{p}$ in the case $d_{0}=d-1$. While we could replace $d_{0}$ in the following proof by the explicit value $d-1$, we prefer to use the former notation to showcase how similar arguments are used for the proof of Theorem 1.2 for $d_{0}>d-1$.
Theorem 3.1.
For a $p$-concentric sphere collection $S_{p}$ as outlined above, with $d_{0}=\dim_{B}S_{0}=d-1$, we have
$\dim_{B}S_{p}=\max\left\{\frac{d}{1+p},\,d-1\right\}.$
Proof.
Fix positive $d_{1},d_{2}<d$ close to $d_{0}$ such that $d_{1}<d_{0}<d_{2}$. Due to $\dim_{B}S_{0}=d_{0}$ and (2.1), there is $r_{0}=r_{0}(d_{1},d_{2})>0$ such that
(3.4) $r^{-d_{1}}\leq N_{r}(S_{0})\leq r^{-d_{2}},$
for all $r\in(0,r_{0})$. Moreover, since $f$ is a homeomorphism, the topological dimension of $B_{0}=f(B(0,1))$ is equal to $d$. It follows that $\dim_{B}B_{0}=d$. Hence, for any tiny $\epsilon>0$, there is $r_{0}^{\prime}=r_{0}^{\prime}(\epsilon)>0$ such that
(3.5) $r^{-d+\epsilon}\leq N_{r}(B_{0})\leq r^{-d-\epsilon},$
for all $r\in(0,r_{0}^{\prime})$.
Fix an arbitrary tiny $\epsilon>0$, and an arbitrary $r\in(0,\min\{r_{0},r_{0}^{\prime}\})$, so that (3.4), (3.5) hold, and let $n_{r}$ be the unique positive integer satisfying
(3.6) $n_{r}^{-p}-(n_{r}+1)^{-p}\leq\frac{r}{c_{2}d^{1/2}}<(n_{r}-1)^{-p}-n_{r}^{-p}.$
In arguments that follow, we make a few more assumptions on how small $r$ is selected to be, which only involve uniform constants, such as $p$, $c_{1}$, $c_{2}$, $d$, and the uniform comparability constant for the similarity ratios of $g_{j}$ from (3.3). This ensures there is no loss of generality in these further assumptions on the size of $r$, while it eases the notation and reduces the number of technical inequalities.
By the choice of $n_{r}$ in (3.6), we can split $S_{p}$ into two disjoint sets, i.e., $S_{p}=S_{p}^{+}\cup S_{p}^{-}$, where
$S_{p}^{+}:=\bigcup_{n>n_{r}}S_{n}$
$S_{p}^{-}:=\bigcup_{n\leq n_{r}}S_{n},$
and consider $N_{r}(S_{p}^{+})$ and $N_{r}(S_{p}^{-})$ separately. The reason is that $r$-cubes that intersect $S_{p}^{+}$ are essentially disjoint with those that intersect $S_{p}^{-}$ by (3.2) and (3.6), hence resulting in $N_{r}(S_{p})=N_{r}(S_{p}^{+})+N_{r}(S_{p}^{-})$. Moreover, by (3.2) and the right-hand side of (3.6), the collection of $r$-cubes intersecting some $S_{n}\subset S_{p}^{-}$ for $n\leq n_{r}$ and the collection of $r$-cubes intersecting some $S_{m}$ for $n\neq m\leq n_{r}$ are essentially disjoint collections, implying that $N_{r}(S_{p}^{-})=\sum_{n=1}^{n_{r}}N_{r}(S_{n})$. Moreover, by the left-hand side of (3.6) and by (3.1), (3.3), $r$ is large enough for each cube in the $r$-mesh intersecting $B(0,c_{1}n_{r}^{-p})$ to overlap with more than one of the spheres in $S^{+}$, resulting in $N_{r}(S^{+})\simeq N_{r}(g_{n_{r}}(B_{0}))$, where the comparability constant is independent of $r$. Therefore, we have
(3.7) $N_{r}\left(S_{p}\right)\simeq N_{r}(g_{n_{r}}(B_{0}))+\sum_{j=1}^{n_{r}}N_{r}\left(S_{j}\right)$
Note that $n_{r}$ increases as $r$ decreases. Hence, for small enough $r$ (in a way that depends only on $p,c_{1},c_{2},d$), the relation (3.6) actually implies
(3.8) $n_{r}\simeq r^{-\frac{1}{p+1}},$
where the comparability constant depends only on $p,c_{1},c_{2},d$. The similarity ratio of $g_{n_{r}}$ is comparable to $n_{r}^{-p}$. Since $n_{r}^{p}r\simeq n_{r}^{-1}$ by (3.8), without loss of generality, we can assume that $r$ is small enough, in a way solely dependent on $p$ and the comparability constant in the former relation, so that $n_{r}^{p}r$ is smaller than $\min\{r_{0},r_{0}^{\prime}\}$. This ensures that (2.2) can be applied for the fixed $r$, which along with (3.5) and (3.8) results in
(3.9) $r^{-\frac{d-\epsilon}{1+p}}\simeq\left(\frac{1}{n_{r}^{p}r}\right)^{d-\epsilon}\lesssim N_{r}(g_{n_{r}}(B_{0}))\lesssim\left(\frac{1}{n_{r}^{p}r}\right)^{d+\epsilon}\simeq r^{-\frac{d+\epsilon}{1+p}},$
for the fixed tiny $\epsilon>0$.
Since for any $j\in\{1,\dots,n_{r}\}$ the similarity ratio of $g_{j}$ is comparable to $j^{-p}$, and noting that $j^{p}r\leq n_{r}^{p}r<\min\{r_{0},r_{0}^{\prime}\}$ and $S_{j}=g_{j}(S_{0})$, we similarly have by (2.2) that
(3.10) $\left(\frac{r}{j^{-p}}\right)^{-d_{1}}\lesssim N_{r}(g_{j}(S_{0}))\lesssim\left(\frac{r}{j^{-p}}\right)^{-d_{2}}.$
Combining (3.7), (3.9) and (3.10) results in
(3.11) $r^{-\frac{d-\epsilon}{1+p}}+\sum_{j=1}^{n_{r}}\frac{j^{-pd_{1}}}{r^{d_{1}}}\lesssim N_{r}(S_{p})\lesssim r^{-\frac{d+\epsilon}{1+p}}+\sum_{j=1}^{n_{r}}\frac{j^{-pd_{2}}}{r^{d_{2}}}.$
Suppose that $pd_{1}<1$. Then by standard partial sum estimates and (3.8) we have
(3.12) $\sum_{j=1}^{n_{r}}\frac{j^{-pd_{1}}}{r^{d_{1}}}\simeq n_{r}^{1-pd_{1}}r^{-d_{1}}\simeq r^{-\frac{1-pd_{1}}{p+1}-d_{1}}=r^{\frac{-1-d_{1}}{p+1}}.$
Similarly, if $pd_{1}>1$, we have
(3.13) $\sum_{j=1}^{n_{r}}\frac{j^{-pd_{1}}}{r^{d_{1}}}\simeq\frac{1}{pd_{1}-1}r^{-d_{1}}\simeq r^{-d_{1}}.$
A similar case study on the product $pd_{2}$ shows that for $pd_{2}<1$ we have
(3.14) $\sum_{j=1}^{n_{r}}\frac{j^{-pd_{2}}}{r^{d_{2}}}\simeq n_{r}^{1-pd_{2}}r^{-d_{2}}\simeq r^{-\frac{1-pd_{2}}{p+1}-d_{2}}=r^{\frac{-1-d_{2}}{p+1}},$
and for $pd_{2}>1$ we have
(3.15) $\sum_{j=1}^{n_{r}}\frac{j^{-pd_{2}}}{r^{d_{2}}}\simeq\frac{1}{pd_{2}-1}r^{-d_{2}}\simeq r^{-d_{2}}.$
We now determine $\dim_{B}S_{p}$ based on the value of the product $pd_{0}$. Suppose $pd_{0}<1$, then $d_{2}$ can be chosen close enough to $d_{0}$ to ensure $pd_{1}<pd_{2}<1$. Thus, by (3.11), (3.12), (3.14) we have
$r^{-\frac{d-\epsilon}{1+p}}+r^{\frac{-1-d_{1}}{p+1}}\lesssim N_{r}(S_{p})\lesssim r^{-\frac{d+\epsilon}{1+p}}+r^{\frac{-1-d_{2}}{p+1}}.$
Note that $d_{2}>d-1$, so $\epsilon$ can be chosen small enough so that $d_{2}+1>d+\epsilon$. Since $r$ was arbitrary, the above relation implies that
$\frac{1+d_{1}}{p+1}\leq\dim_{B}S_{p}\leq\frac{1+d_{2}}{p+1}.$
Recall that $d_{1}$, $d_{2}$ are arbitrary and as close to $d_{0}$ as necessary, which combined with the above inequality is enough to show that in the case $pd_{0}<1$, we have
$\dim_{B}S_{p}=\frac{1+d_{0}}{p+1}=\frac{d}{p+1}.$
On the other hand, if $pd_{0}>1$, we similarly make sure $pd_{2}>pd_{1}>1$ by choice of $d_{1},d_{2}$. Note that by $d_{2}>d-1$, the value $\epsilon$ can be chosen small enough so that $pd_{2}>1$ implies that $d+\epsilon<pd_{2}+d_{2}$, which gives
$\frac{d+\epsilon}{p+1}<d_{2}.$
Using a similar argument to that of the previous case, and the above inequality, along with (3.11), (3.13), (3.15), we can show that
$d_{1}\leq\dim_{B}S_{p}\leq d_{2},$
which is enough to conclude that $\dim_{B}S_{p}=d_{0}$ as needed.
The case $pd_{0}=1$ is treated similarly, using (3.11), (3.12), (3.15), resulting in $\dim_{B}S_{p}=d_{0}=d-1$, due to the fact that
$\frac{1+d_{0}}{1+p}=\frac{1+d_{0}}{1+1/d_{0}}=d_{0}.$
All cases have been addressed and the proof is complete.
Remark 3.2.
It might be tempting to consider $r\simeq n_{r}^{-p}$, so that all but finitely many of the spheres $S_{j}\subset S_{p}$ are covered by the $2^{d}$ $r$-cubes with vertices at 00. However, this choice of $n_{r}$ would hinder any following covering arguments, and especially the crucial equation $N_{r}(S_{p}^{-})=\sum_{j=1}^{n_{r}}N_{r}(S_{j})$ would no longer be true.
The techniques employed in the proof of Theorem 3.1 are in fact very similar to those needed to determine the Assouad spectrum of $S_{p}$. By adjusting the right quantitative estimates, we can modify accordingly the arguments within the previous proof, so that they apply in the context of two small scales, $r$ and $r^{\theta}$, as in the definition of the $\theta$-Assouad spectrum.
Theorem 3.3.
[…]
(3.16) $\dim_{A}^{\theta}S_{p}=\min\left\{\frac{d}{(1+p)(1-\theta)},\,d\right\},$
for all $\theta\in(0,1)$. On the other hand, if $p(d-1)>1$, then
(3.17) $\dim_{A}^{\theta}S_{p}=\min\left\{d-1+\frac{\theta}{p(1-\theta)},\,d\right\},$
for all $\theta\in(0,1)$.
Proof.
Recall that we assume $d_{0}=d-1=\dim_{A}S_{0}$. Note that for any $\theta\geq p/(p+1)$, both (3.16) and (3.17) result in $\dim_{A}^{\theta}S_{p}=d$. Hence, by continuity of the Assouad spectrum as a function of $\theta$, it is enough to show (3.16) and (3.17) for all $\theta<p/(p+1)$.
Fix $\theta\in(0,p/(p+1))$ and positive $d_{1},d_{2}$ close to $d_{0}$, such that $d_{1}<d_{0}<d_{2}$. Moreover, fix $r\in(0,1)$, $\epsilon>0$ as small as necessary for the following arguments, in a way that depends only on uniform constants of $S_{p}$ and $\mathbb{R}^{d}$, similarly to the proof of Theorem 3.1. Denote by $n_{r}\in\mathbb{N}$ the unique positive integer satisfying
(3.18) $n_{r}^{-p}-(n_{r}+1)^{-p}\leq\frac{r}{c_{2}d^{1/2}}<(n_{r}-1)^{-p}-n_{r}^{-p},$
and by $m_{r}\in\mathbb{N}$ the unique positive integer satisfying
(3.19) $m_{r}^{-p}\leq\frac{r^{\theta}}{c_{1}}<(m_{r}-1)^{-p}.$
Since $r$ was fixed to be small enough, the integers $n_{r}$, $m_{r}$ are large enough, so that by (3.18) we have
(3.20) $n_{r}\simeq r^{-\frac{1}{p+1}},$
and by (3.19) we have
(3.21) $m_{r}\simeq r^{-\frac{\theta}{p}}.$
Due to the choice $\theta<p/(p+1)$, by (3.18) and (3.19) we also have that $m_{r}\leq n_{r}$. Note that the Assouad spectrum is stable under taking the closure of a set (see [22, p.18, 49]), which ensures that we may estimate the Assouad spectrum of the closure of $S_{p}$, or simply assume that $0\in S_{p}$ without loss of generality. Moreover, despite the fact that the definition of the Assouad spectrum requires to check balls centered at all points of the given set, we focus on $B(0,r^{\theta})$ instead, for reasons that become clear later.
Arguing similarly to the proof of Theorem 3.1 and splitting $B(0,r^{\theta})\cap S_{p}$ into two disjoint sets, we have
$N_{r}(B(0,r^{\theta})\cap S_{p})\simeq N_{r}(g_{n_{r}}(B_{0}))+\sum_{j=m_{r}}^{n_{r}}N_{r}(S_{j}),$
which by (2.2) and the choices of $\epsilon,d_{1},d_{2}$ implies that
$\sum_{j=m_{r}}^{n_{r}}\frac{j^{-pd_{1}}}{r^{d_{1}}}\lesssim N_{r}(B(0,r^{\theta})\cap S_{p})\lesssim\left(\frac{n_{r}^{-p}}{r}\right)^{d+\epsilon}+\sum_{j=m_{r}}^{n_{r}}\frac{j^{-pd_{2}}}{r^{d_{2}}}.$
Using (3.20) and standard partial summation techniques on the above relation results in
(3.22) $r^{-d_{1}}\max\{m_{r}^{1-pd_{1}},n_{r}^{1-pd_{1}}\}\lesssim N_{r}(B(0,r^{\theta})\cap S_{p})\lesssim r^{\left(\frac{p}{p+1}-1\right)(d+\epsilon)}+r^{-d_{2}}\max\{m_{r}^{1-pd_{2}},n_{r}^{1-pd_{2}}\}.$
The rest of the proof is a case study on the product $pd_{0}$, similarly to the proof of Theorem 3.1, while also accounting for the different values of $\theta$ that could potentially determine which power of $r$ is dominant.
Suppose $pd_{0}<1$. In this case, we only need to find a sequence of balls centered at $S_{p}$ and bound the corresponding covering number from below, which yields a lower bound on $\dim_{A}^{\theta}S_{p}$. This is because the desired upper bound on $\dim_{A}^{\theta}S_{p}$ follows trivially by Theorem 3.1 and Lemma 2.1. Hence, we can focus on balls centered at 00 and the lower bound of (3.22). Namely, by (3.22) and $pd_{1}<1$ we have
$r^{-d_{1}}n_{r}^{1-pd_{1}}\lesssim N_{r}(B(0,r^{\theta})\cap S_{p}),$
which by (3.20) implies
$N_{r}(B(0,r^{\theta})\cap S_{p})\gtrsim r^{-d_{1}-\frac{1-pd_{1}}{p+1}}=(r^{\theta-1})^{\frac{-d_{1}(p+1)-(1-pd_{1})}{(\theta-1)(p+1)}}=(r^{\theta-1})^{\frac{d_{1}+1}{(1-\theta)(p+1)}}.$
Thus, the above shows that
$\dim_{A}^{\theta}S_{p}\geq\frac{d_{1}+1}{(1-\theta)(p+1)},$
for arbitrary $d_{1}$ close to $d_{0}$, which is enough to complete the proof of (3.16).
Suppose that $pd_{0}>1$. Note at this point that if $x\in S_{p}$ with $|x|\leq 2r^{\theta}$, then
(3.23) $N_{r}(B(x,r^{\theta})\cap S_{p})\leq N_{r}(B(0,10r^{\theta})\cap S_{p})\lesssim N_{r}(B(0,r^{\theta})\cap S_{p}),$
where the comparability constant only depends on the doubling constant of the space $\mathbb{R}^{d}$ (see for instance [22, Section 13.1]). Hence, it remains to study the covering number $N_{r}(B(0,r^{\theta})\cap S_{p})$, and $N_{r}(B(x,r^{\theta})\cap S_{p})$ for arbitrary $x\in S_{p}$ with $|x|>2r^{\theta}$. In the latter case, we observe that the individual spheres $S_{j}$ that intersect $B(x,r^{\theta})$ are not more than $m_{r}$ in number, due to (3.3) and (3.19) (see also [23, p. 3263] for a similar argument). Therefore, using the fact that $\dim_{A}S_{0}=d_{0}=\dim_{B}S_{0}$, which implies that $\dim_{A}^{\theta}S_{0}=d_{0}$ by (2.4), and a similarity argument similar to (2.2), we have
$N_{r}(B(x,r^{\theta})\cap S_{p})\lesssim m_{r}(r^{\theta-1})^{d_{2}}.$
By (3.21), the above implies that
(3.24) $N_{r}(B(x,r^{\theta})\cap S_{p})\lesssim r^{-\frac{\theta}{p}}(r^{\theta-1})^{d_{2}}=(r^{\theta-1})^{d_{2}+\frac{\theta}{p(1-\theta)}}.$
Hence, due to (3.23) and the above inequality, it is enough to focus on achieving lower and upper bounds for $N_{r}(B(0,r^{\theta})$ in the rest of the proof. The relation $pd_{0}>1$ implies for (3.22) that
$r^{-d_{1}}m_{r}^{1-pd_{1}}\lesssim N_{r}(B(0,r^{\theta})\cap S_{p})\lesssim r^{\left(\frac{p}{p+1}-1\right)(d+\epsilon)}+r^{-d_{2}}m_{r}^{1-pd_{2}}.$
Using (3.21) in the above relation results in
$r^{-d_{1}}r^{-\frac{\theta(1-pd_{1})}{p}}\lesssim N_{r}(B(0,r^{\theta})\cap S_{p})\lesssim r^{\frac{-(d+\epsilon)}{p+1}}+r^{-d_{2}}r^{-\frac{\theta(1-pd_{2})}{p}}.$
Note that
$-\frac{d_{1}}{\theta-1}-\frac{\theta(1-pd_{1})}{p(\theta-1)}=d_{1}+\frac{\theta}{p(1-\theta)},$
and similarly for $d_{2}$ instead of $d_{1}$, which when applied to the previous inequality yield
(3.25) $(r^{\theta-1})^{d_{1}+\frac{\theta}{p(1-\theta)}}\lesssim N_{r}(B(0,r^{\theta})\cap S_{p})\lesssim(r^{\theta-1})^{\frac{d+\epsilon}{(p+1)(1-\theta)}}+(r^{\theta-1})^{d_{2}+\frac{\theta}{p(1-\theta)}}.$
Regarding the dominant term in the upper bound of (3.25), note that
$d_{0}+\frac{\theta}{p(1-\theta)}>\frac{d}{(p+1)(1-\theta)}$
if, and only if
$\theta<\frac{p(pd_{0}-d_{0}-d)}{(p+1)(pd_{0}-1)}=\frac{p}{p+1}.$
Hence, by choice of $\theta$, and by taking $\epsilon$ to be small enough, we have
$d_{2}+\frac{\theta}{p(1-\theta)}>\frac{d+\epsilon}{(p+1)(1-\theta)},$
which if applied to (3.25) yields
$(r^{\theta-1})^{d_{1}+\frac{\theta}{p(1-\theta)}}\lesssim N_{r}(B(0,r^{\theta})\cap S_{p})\lesssim(r^{\theta-1})^{d_{2}+\frac{\theta}{p(1-\theta)}}.$
As a result, by the above and by (3.23),(3.24), we have shown that
$d_{1}+\frac{\theta}{p(1-\theta)}\leq\dim_{A}^{\theta}S_{p}\leq d_{2}+\frac{\theta}{p(1-\theta)},$
for $d_{1}$ and $d_{2}$ arbitrarily close to $d_{0}$, which is enough to complete the proof in this case as well.
The case $pd_{0}=1$ is treated similarly to the above cases. Notice that in this case we have $pd_{1}<1$ and $1<pd_{2}$, which yield different dominant terms on the two sides of (3.22), but the arguments are identical for each side to the corresponding cases; namely, the lower bound is treated using $pd_{1}<1$ as in the case $pd_{0}<1$, and the upper bound is treated using $pd_{2}>1$ as in the case $pd_{0}>1$. This is enough to complete the proof ∎
Remark 3.4.
We emphasize that in the case $p(d-1)>1$, the Assouad spectrum is not given by the upper bound in Lemma 2.1 for $0<\theta<\frac{p}{1+p}$. This provides us with a family of such examples in every Euclidean space $\mathbb{R}^{d}$.
Theorem 3.3 in fact implies what the box dimension of $S_{p}$ is equal to as well, by letting $\theta\rightarrow 0$ (see Section 2). We note that for $p>(d-1)^{-1}$, the box dimension of $S_{p}$ is trivial, in the sense that $\dim_{B}S_{p}=d-1$. On the other hand, the Assouad spectrum provides non-trivial information on $S_{p}$, even in the case where the box dimension fails to do so. Moreover, (3.16), (3.17) yield that the Assouad dimension of $S_{p}$ is equal to $d$, which also fails to provide additional geometric information compared to the Assouad spectrum. This is an immediate implication of (2.4).
Corollary 3.5.
For all $p>0$, we have $\dim_{A}S_{p}=\dim_{\mathrm{qA}}S_{p}=d$, where $S_{p}$ is as in Theorem 3.3.
We finish this section by showing that a topological sphere of large box dimension cannot generate a polynomially concentric sphere collection.
Proof of Theorem 1.2.
Assume towards contradiction that $S_{p}$ is a $p$-polynomially concentric sphere collection, as described in the beginning of Section 3, but with $d_{0}=\dim_{B}S_{0}>d-1$. Note that we do not require $S_{0}$ to satisfy $\dim_{A}S_{0}=d_{0}$. We arrive at a contradiction by providing an absurd lower bound on the Assouad spectrum of $S_{p}$ for certain $\theta\in(0,1)$.
Fix $d_{1}<d_{0}$ close to $d_{0}$, such that $d_{1}>d-1$, and $\theta\in(0,\frac{p}{p+1})$. Moreover, similarly to the proof of Theorem 3.3, fix small $r\in(0,1)$ and integers $n_{r}$, $m_{r}$ so that (3.18), (3.19) are satisfied. Choosing $r$ small enough ensures that $n_{r}\simeq r^{-\frac{1}{p+1}}$ and $m_{r}\simeq r^{-\frac{\theta}{p}}$, which by choice of $\theta<p/(p+1)$ also ensures that $m_{r}\leq n_{r}$. Arguing exactly as in the proof of Theorem 3.3 yields
(3.26) $N_{r}(B(0,r^{\theta})\cap S_{p})\gtrsim\sum_{j=m_{r}}^{n_{r}}\frac{j^{-pd_{1}}}{r^{d_{1}}}\gtrsim r^{-d_{1}}\max\{m_{r}^{1-pd_{1}},n_{r}^{1-pd_{1}}\}.$
Suppose $pd_{0}>1$. Then $d_{1}$ can also be chosen so that $pd_{1}>1$, which implies by (3.26) that
$N_{r}(B(0,r^{\theta})\cap S_{p})\gtrsim(r^{\theta-1})^{d_{1}+\frac{\theta}{p(1-\theta)}},$
exactly as in the proof of Theorem 3.3. Since $r$, $\theta$ are arbitrary, this implies that
$\dim_{A}^{\theta}S_{p}\geq d_{1}+\frac{\theta}{p(1-\theta)},$
which also yields
(3.27) $\dim_{A}^{\theta}S_{p}\geq d_{0}+\frac{\theta}{p(1-\theta)},$
because $d_{1}<d_{0}$ can be arbitrarily close to $d_{0}$. However, note that $d_{0}+\frac{\theta}{p(1-\theta)}<d$ if, and only if, $\theta<\frac{p(d-d_{0})}{p(d-d_{0})+1}$. Since $\frac{p(d-d_{0})}{p(d-d_{0})+1}<\frac{p}{p+1}$, due to $d_{0}>d-1$, if we pick
$\frac{p(d-d_{0})}{p(d-d_{0})+1}<\theta<\frac{p}{p+1},$
then (3.27) implies that $\dim_{A}^{\theta}S_{p}>d,$ which is a contradiction.
On the other hand, if $pd_{0}\leq 1$, we have $pd_{1}<1$, which implies by (3.26) that
$N_{r}(B(0,r^{\theta})\cap S_{p})\gtrsim(r^{\theta-1})^{\frac{d_{1}+1}{(1-\theta)(p+1)}}.$
This similarly yields
$\dim_{A}^{\theta}S_{p}\geq\frac{d_{0}+1}{(1-\theta)(p+1)},$
which by choosing
$\frac{d(p+1)-(d_{0}+1)}{d(p+1)}<\theta<\frac{p}{p+1},$
provides the contradicting estimate $\dim_{A}^{\theta}S_{p}>d$ for this case as well, completing the proof. ∎
Remark 3.6.
For $p<d^{-1}(1+d_{0}-d)$, if we let $d_{1}<d_{0}$ arbitrarily close to $d_{0}$ and follow the argument of the proof of Theorem 3.1 just for the lower bound of $N_{r}(S_{p})$, we get that $\dim_{B}S_{p}\geq\frac{d_{0}+1}{1+p}>d$, by choice of $p$, which is a contradiction. Hence, the box dimension of $S_{p}$ can also trace the impossibility of such a concentric collection, but only for certain values of $p$. Therefore, the investigation of the Assouad spectrum is truly necessary.
4. Quasiconformal classification of spiral shells
As mentioned in the Introduction, quasiconformal maps are a natural generalization of conformal maps in higher dimensions. There are multiple ways to express the intuition behind quasiconformality, i.e., the property that infinitesimal balls are mapped onto infinitesimal ellipsoids whose eccentricity is globally bounded by a constant. We state the analytic definition of quasiconformal maps, which relates directly to the dimension distortion result of interest (see [41] for a thorough exposition). A homeomorphism $f:\Omega\to\Omega^{\prime}$ between domains in $\mathbb{R}^{d}$, $d\geq 2$, is said to be $K$-quasiconformal, for some $K\geq 1$, if $f$ lies in the local Sobolev space $W^{1,d}_{\text{loc}}(\Omega:\mathbb{R}^{d})$, and the inequality
$|Df|^{d}\leq K\det Df$
holds almost everywhere $\Omega$ (with respect to the $d$-Lebesgue measure). Note that $Df$ denotes the (a.e. defined) differential matrix, and $|\mathbf{A}|=\max\{|\mathbf{A}(\mathbf{v})|:|\mathbf{v}|=1\}$ denotes the operator norm of a matrix $\mathbf{A}$. Recall that the local Sobolev space $W^{1,d}_{\text{loc}}(\Omega:\mathbb{R}^{d})$ is the space of maps defined in $\Omega$ with first-order weak derivatives in $L^{p}_{\text{loc}}(\Omega:\mathbb{R}^{d})$ (see [26] for more details).
There are many dimension distortion results, for different analytic and geometric mapping classes related to quasiconformal maps, and regarding different dimension notions. See [25], [30], [4], [5], [16], [15], [14] for a non-exhaustive list. In fact, the following theorem proved by Tyson and the author (see [16, Theorem 1.3]) is crucial for the proof of Theorem 1.3.
Theorem 4.1.
Let $F:\Omega\to\Omega^{\prime}$ be a $K$-quasiconformal map between domains $\Omega,\,\Omega^{\prime}\subset\mathbb{R}^{d}$. For $t>0$ , set $\theta(t)=1/(t+1)$. If $E$ is a non-empty bounded set with closure $\overline{E}\subset\Omega$, we have
(4.1) $\left(1-\frac{d}{s}\right)\left(\frac{1}{\dim_{A,reg}^{\theta(t/K)}E}-\frac{1}{d}\right)\leq\frac{1}{\dim_{A,reg}^{\theta(t)}F(E)}-\frac{1}{d}\leq\left(1-\frac{d}{s}\right)^{-1}\left(\frac{1}{\dim_{A,reg}^{\theta(Kt)}E}-\frac{1}{d}\right),$
for a fixed $s=s(d,K)>d$ that only depends on the space $\mathbb{R}^{d}$ and the constant $K$ of the map $F$.
We have all the tools necessary to fully classify the polynomial spiral shells up to quasiconformal equivalence.
Proof of Theorem 1.3.
One direction is immediate, as there exists a map $F$ with $F(\mathcal{S}_{p})=\mathcal{S}_{q}$, which is $K$-quasiconformal for any $K\geq p/q$. Namely, the radial stretch map $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with $F(0)=0$ and
$F(x)=|x|^{q/p-1}x,$
for all $x\in\mathbb{R}^{d}\setminus\{0\}$, is $(p/q)$-quasiconformal (see [41, Section 16.2]), which implies that $F$ is $K$-quasiconformal for all $K\geq p/q$. Moreover, it is elementary to show that $F$ maps $\mathcal{S}_{p}$ onto $\mathcal{S}_{q}$, as required.
For the other implication, assume towards contradiction that $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a $K$-quasiconformal map with $K<p/q$ and $F(\mathcal{S}_{p})=\mathcal{S}_{q}$. We aim to calculate the regularized Assouad spectrum of the shells $\mathcal{S}_{p},\mathcal{S}_{q}$, and use Theorem 4.1 to arrive at a contradiction. Set $S_{p}:=\cup_{j=1}^{\infty}S(0,j^{-p})$ and $S_{q}:=\cup_{j=1}^{\infty}S(0,j^{-q})$. The sets $S_{p}$, $S_{q}$ are polynomially concentric sphere collections centered at 00 and generated by the unit sphere $S(0,1)$ in $\mathbb{R}^{3}$. Hence, by Theorem 3.3 we have that
[…]
and similarly for $\dim_{A}^{\theta}S_{q}$.
Note that a bi-Lipschitz map $h_{1}$ can be constructed in order to map $\mathcal{S}_{p}^{\text{u}}:=\mathcal{S}_{p}\cap\{(x_{1},x_{2},x_{3}):\,x_{3}\geq 0\}$ onto $S_{p}^{\text{u}}:=S_{p}\cap\{(x_{1},x_{2},x_{3}):\,x_{3}\geq 0\}$ (see [12, 23] for similar arguments). Similarly, a different bi-Lipschitz map $h_{2}$ can be constructed in order to map $\mathcal{S}_{p}^{\ell}:=\mathcal{S}_{p}\cap\{(x_{1},x_{2},x_{3}):\,x_{3}\leq 0\}$ onto $S_{p}^{\ell}:=S_{p}\cap\{(x_{1},x_{2},x_{3}):\,x_{3}\leq 0\}$. Using the finite and bi-Lipschitz stability of the Assouad spectrum (see [22, p. 18, 49], we have
$\dim_{A}^{\theta}\mathcal{S}_{p}=\max\{\dim_{A}^{\theta}\mathcal{S}_{p}^{\text{u}},\dim_{A}^{\theta}\mathcal{S}_{p}^{\ell}\}=\max\{\dim_{A}^{\theta}S_{p}^{\text{u}},\dim_{A}^{\theta}S_{p}^{\ell}\}=\dim_{A}^{\theta}S_{p},$
for all $\theta\in(0,1)$. Similarly we can show that $\dim_{A}^{\theta}\mathcal{S}_{q}=\dim_{A}^{\theta}S_{q}$. Hence, by the above and (4.2) we can fully determine the Assouad spectrum of $\mathcal{S}_{p}$ and $\mathcal{S}_{q}$. Moreover, since the expressions in (4.2) are monotone, by the relation of the Assouad spectrum with the regularized Assouad spectrum in (2.3), we have
[…]
and similarly for $\dim_{A,reg}^{\theta}\mathcal{S}_{q}$.
We have all the necessary dimension-theoretic information in order to apply Theorem 4.1. In fact, we mostly use the relation between the phase transitions $\theta_{p}=p/(p+1)$ and $\theta_{q}=q/(q+1)$ of the two shells, i.e., the values where the regularized Assouad spectrum reaches the (quasi-)Assouad dimension. Set $t=1/q$. Then $\theta(t)=1/(1+t)=q/(1+q)$, and so $\dim_{A}^{\theta(t)}(S_{q})=3$ by (4.3). On the other hand,
$\theta(t/K)=\frac{K}{K+t}=\frac{qK}{qK+1}<\frac{p}{p+1},$
due to the fact that $K<p/q$. Therefore, $\dim_{A}^{\theta(t/K)}(S_{p})<3$. This leads to a contradiction due to Theorem 4.1 and, in particular, due to the left-hand side of (4.1). Therefore, such a $K$-quasiconformal map does not exist. ∎
Remark 4.2.
A closer analysis of the above proof indicates that Theorem 1.3 could be generalized to non-global quasiconformal maps. More specifically, the result is still be true if the map $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ in the statement of the theorem is replaced by a map $F:\Omega\rightarrow\Omega^{\prime}$ between domains of $\mathbb{R}^{d}$ such that $\mathcal{S}_{p}\cup\{0\}\subset\Omega$ and $\mathcal{S}_{q}\cup\{0\}\subset\Omega^{\prime}$.
We emphasize that the Assouad spectrum estimates are crucial in the proof of Theorem 1.3. Indeed, in the case $1/2\leq q\leq p$, all other dimension notions do not provide non-trivial information, since $\dim_{H}\mathcal{S}_{p}=\dim_{B}\mathcal{S}_{p}=2<3=\dim_{A}\mathcal{S}_{p}$, where $\dim_{H}$ denotes the Hausdorff dimension (see [18]), and similarly for $\mathcal{S}_{q}$. The available dimension distortion results for the Hausdorff, box, and Assouad dimension under quasiconformal maps are not enough to classify such polynomial shells. See [16] for a more extensive discussion on this phenomenon in the case of polynomial spirals.
5. Final Remarks
Suppose $S_{0}=f(S(0,1))\subset\mathbb{R}^{d}$ is a topological sphere, where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a homeomorphism, and $x_{0}$ is a point lying in the interior of the topological ball $f(B(0,1))$. We address the existence of a concentric sphere collection generated by $S_{0}$, regardless of the value of $\dim_{B}S_{0}\in[d-1,d]$. Since $f(B(0,1))$ is an open set, there is a maximal open ball lying in $f(B(0,1))$, centered at $x_{0}$ and of radius $r_{1}>0$. Set $g_{1}$ to be a similarity such that $x_{0}\in g_{1}(B(0,1))\subset B(x_{0},r_{1})$. Repeating the same argument for $g_{1}(B(0,1))$ and iterating infinitely many times, we construct a sequence of similarities $\{g_{n}\}_{n\in\mathbb{N}}$ that contracts $S_{0}$ into the point $x_{0}$, i.e., a concentric sphere collection centered at $x_{0}$ and generated by $S_{0}$. However, by Theorem 1.2, we have established that the similarity ratios of $g_{n}$ cannot be comparable to any sequence of the form $n^{-p}$, for any $p>0$, unless $\dim_{B}S_{0}=d-1$. Thus, in the case $\dim_{B}S_{0}\in(d-1,d)$, the sequence $a_{n}$ in the constructed example has to converge to 00 faster than $n^{-p}$ does. Note that in such situations, the convergence is typically too fast for any dimension notion to trace “fractalness” in the collection. For instance, following the method of the proof of Theorem 3.3, it can be shown that a concentric sphere collection with sequence $a_{n}=e^{-n}$ has Assouad spectrum (and, similarly, Assouad dimension) equal to $\dim_{B}S_{0}$. A natural question motivated by this discussion is to try and fully characterize the topological spheres $S_{0}$ from which a polynomially concentric collection can be generated. Given a topological sphere $S_{0}\subset\mathbb{R}^{d}$, Theorem 1.2 shows that $\dim_{B}S_{0}=d-1$ is a necessary condition, but is it sufficient? Namely, if $\dim_{B}S_{0}=d-1$, can we polynomially contract $S_{0}$ into a point using similarities? If the answer to this question is positive in general, are there specific restrictions imposed on the degree $p$ of the polynomial rate $n^{-p}$, based on properties of $S_{0}$?
It should be emphasised that the case $p>(d-1)^{-1}$ in Theorem 3.3 is another one of the few examples of sets $E\subset\mathbb{R}^{d}$, where $\dim_{A}^{\theta}E$ is not constant and is not equal to $\dim_{B}E/(1-\theta)$, for $\theta<1-\dim_{B}E/d$ (see Lemma 2.1). Many of the first such examples mostly appeared in $\mathbb{R}^{2}$, such as the polynomial spirals studied by Fraser in [23], and the elliptical polynomial spirals studied by Falconer-Fraser-Burrell in [12] . In fact, the examples studied in the latter paper stand out for one more reason; for fixed $p,q$, the Assouad spectrum of elliptical spirals $\mathcal{S}_{p,q}$ exhibits two distinct phase transitions, i.e., $\dim_{A}^{\theta}\mathcal{S}_{p,q}$ has three different branches as a function, unlike $\dim_{A}^{\theta}\mathcal{S}_{p}$ which only has two (see [12, Theorem 2.6]). This motivates the study of another class in $\mathbb{R}^{d}$, namely that of concentric ellipsoid collections. In particular, this class is defined similarly to the concentric sphere collections in Definition 1.1, but with the second condition (ii) replaced by $\operatorname{dist}(S_{n},S_{n+1})=c_{2}b_{n}$, for some strictly decreasing sequence $\{b_{n}\}_{n\in\mathbb{N}}$ that converges to 00 at a different rate than $\{a_{n}\}_{n\in\mathbb{N}}$. We expect the dimension study of these objects to be quite interesting, potentially providing us with more examples of sets with Assouad spectrum of two distinct phase transition values.
Recall that a homeomorphism $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is $(\alpha,\beta)$-bi-Hölder continuous on a non-empty set $E$, with Hölder exponents $0<\alpha\leq 1\leq\beta<\infty$, if there is a constant $C>0$ such that
$C^{-1}|x-y|^{\beta}\leq|f(x)-f(y)|\leq C|x-y|^{\alpha},$
for all $x,y\in E$. Fraser in [23] suggested the following classification-like direction: given two bounded sets $X,Y\subset\mathbb{R}^{d}$, try to construct a bi-Hölder map $f$ such that $f(X)=Y$. Moreover, determine what is the sharpest possible choice of Hölder exponents for such a map (which depends on properties of $X,Y$), and compare them to the bounds on the exponents from dimension distortion theorems. We have indirectly stepped towards this direction for spiral shells through Theorem 1.3, and the connection between quasiconformal and bi-Hölder maps. Namely, any $K$-quasiconformal map $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is in fact $(1/K,K)$-bi-Hölder continuous on every compact set $E\subset\mathbb{R}^{d}$ (see for instance [28, Theorem 7.7.1]). Hence, in the proof of Theorem 1.3, we provided a $(q/p,p/q)$-bi-Hölder map that maps $\mathcal{S}_{p}\cup\{0\}$ onto $\mathcal{S}_{q}\cup\{0\}$. However, while this map is sharp for the quasiconformal classification of spiral shells, it is not clear whether it is sharp for the bi-Hölder class of maps between such shells.
One could use the bi-Hölder dimension distortion results for the Assouad spectrum (see [22, Section 3.4.3]), and try to come up with bounds on the Hölder exponents of a map that sends $\mathcal{S}_{p}$ onto $\mathcal{S}_{q}$ in terms of $p$ and $q$. For instance, if $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is $(\alpha,\beta)-$bi-Hölder and $E\subset\mathbb{R}^{d}$ is bounded, then
$\frac{\dim_{B}E}{\beta}\leq\dim_{B}f(E)\leq\frac{\dim_{B}E}{\alpha}.$
Therefore, in the case $10<q<p<$ 1, using the the box dimension of $\mathcal{S}_{p},\mathcal{S}_{q}$ and the above inequality, we arrive at the trivial bound $\beta\geq(q+1)/(p+1)$, and the bound $\alpha\leq(q+1)/(p+1)$ for the Hölder exponents. This does not necessarily imply that a $\frac{q+1}{p+1}$-Hölder map $f$ with $f(\mathcal{S}_{p})=\mathcal{S}_{q}$ actually exists. Note that for $0<q<p$ we have
$\frac{q}{p}<\frac{q+1}{p+1},$
so the latter term is indeed a “better” Hölder exponent, in the sense that it leads to a “smoother” map $f$. However, it appears that apart from certain specialized examples, the Hölder dimension distortion estimates are rarely sharp. This was also emphasized in the Hölder unwinding problem for polynomial spirals by Fraser in [23]. As a result, it is not unreasonable to suspect that $q/p$ is the sharpest exponent possible, at least for the classification problem involving $(\alpha,\alpha^{-1})$-bi-Hölder maps between spiral shells $\mathcal{S}_{p},\mathcal{S}_{q}$. In fact, a natural direction to explore is the following; suppose the sharp bi-Hölder exponent bounds in a classification problem are not achieved through knowledge of the dimensions of the sets and the Hölder dimension distortion estimates, are the bounds achieved by the corresponding quasiconformal classification results sharp?